### v1

# Encoder vs Decoder: Comparative Analysis of Encoder and Decoder Language Models on Multilingual NLU Tasks

Abstract
This paper explores the performance of encoder and decoder language models on multilingual Natural Language Understanding (NLU) tasks, with a broad focus on Germanic languages. Building upon the ScandEval benchmark, which initially was restricted to evaluating encoder models, we extend the evaluation framework to include decoder models. We introduce a method for evaluating decoder models on NLU tasks and apply it to the languages Danish, Swedish, Norwegian, Icelandic, Faroese, German, Dutch, and English. Through a series of experiments and analyses, we address key research questions regarding the comparative performance of encoder and decoder models, the impact of NLU task types, and the variation across language resources. Our findings reveal that decoder models can achieve significantly better NLU performance than encoder models, with nuances observed across different tasks and languages. Additionally, we investigate the correlation between decoders and task performance via a UMAP analysis, shedding light on the unique capabilities of decoder and encoder models. This study contributes to a deeper understanding of language model paradigms in NLU tasks and provides valuable insights for model selection and evaluation in multilingual settings.
Encoder vs Decoder: Comparative Analysis of Encoder and Decoder Language Models on Multilingual NLU Tasks
Dan Saattrup Nielsen The Alexandra Institute dan.nielsen@alexandra.dk                        Kenneth Enevoldsen University of Aarhus kenneth.enevoldsen@cas.au.dk                        Peter Schneider-Kamp University of Southern Denmark petersk@imada.sdu.dk
1 Introduction
Language models have attained remarkable Natural Language Understanding (NLU) performance, both with encoder-based architectures like BERT Devlin et al. (2018) and and decoder-based architectures like GPT-3 Brown, Tom and Mann, Benjamin and Ryder, Nick and Subbiah, Melanie and Kaplan, Jared D and Dhariwal, Prafulla and Neelakantan, Arvind and Shyam, Pranav and Sastry, Girish and Askell, Amanda and others (2020). The encoder models have excelled in capturing contextual information for downstream tasks through masked language modeling objectives, while decoder models have shown strong generative capabilities by autoregressively predicting subsequent tokens based on preceding context.
Since the “ChatGPT boom” in 2023, the research community has been increasingly focused on decoder models Zhao et al. (2023) for both Natural Language Generation (NLG) and NLU tasks. However, few studies have systematically compared the performance of encoder and decoder models across a diverse range of NLU tasks, and the studies that exist have primarily focused on English. This leaves a gap in our understanding of how the two language model paradigms perform in multilingual settings across different languages and tasks.
Nielsen (2023) introduced the ScandEval benchmark and evaluated encoder language models on four different natural language understanding tasks in Danish, Swedish, Norwegian (Bokmål and Nynorsk), Icelandic and Faroese. In this paper, we bridge this gap by extending the ScandEval benchmark to encompass the evaluation of decoder models on multilingual NLU tasks, as well as expanding the language resources to include German, Dutch and English.
Our main research question is
Which language model paradigm is better suited for NLU?
We will answer this question with the languages Danish, Swedish, Norwegian, Icelandic, Faroese, German, Dutch and English as a case study. To concretise our main question, we will study the following research questions in this paper:
  1. (Q1)
    Can state-of-the-art decoder models achieve significantly better NLU performance than state-of-the-art finetuned encoder models?
  2. (Q2)
    Does the answer to (Q1) depend on the type of NLU task?
  3. (Q3)
    Does the answer to (Q1) vary along the language resource spectrum, from low- to high-resource?
Our main contributions of this paper are the following:
  1. 1.
    We extend the ScandEval benchmarking framework with few-shot evaluation of decoder models and release this extension open-source.
  2. 2.
    We extend the languages supported by the ScandEval benchmarking framework by German, Dutch and English. Together with Danish, Swedish, Norwegian, Icelandic and Faroese, ScandEval now provides virtually full coverage of the Germanic language group.
  3. 3.
    We evaluate an extensive suite of both encoder and decoder models on NLU tasks in all of the supported languages and publish these on public leaderboards.
  4. 4.
    We give a positive answer to (Q1), showing that decoder models can achieve significantly better NLU performance than encoder models. This depends on the language in question however, giving a partially positive answer to (Q3).
  5. 5.
    We also show that the decoder models are heavily biased towards the question answering task (even models that are not instruction tuned), and a UMAP analysis shows that the performance distribution of decoder models follow a different “path” than encoder models, from the worst to best performing models. This gives a positive answer to (Q2).
2 Related Work
2.1 Comparing Encoder and Decoder Models
There has been a number of studies in recent years comparing encoder models to decoder models. Zhong et al. (2023) compared GPT-3.5-turbo (January 2023 version) to (finetuned versions of) the base- and large versions of BERT Devlin et al. (2018) and RoBERTa Liu et al. (2019) on the English GLUE benchmark Wang et al. (2018). They found that GPT-3.5-turbo was on average on par with the base-sized encoder models, but fell short of the large-sized ones. They also note that despite being on par with the base-sized models, there was a big discrepancy between the models on individual tasks, with GPT-3.5-turbo for instance being better on the inference tasks while being worse on the paraphrase tasks. We note however that they only evaluate the decoder model in a zero-shot setting, and furthermore they only evaluate the models on 25 samples for each class in the development split, leading to a potential lack of robustness in their evaluation.
Wang et al. (2023) compared GPT-3.5-turbo (January 2023 version) to a finetuned version of the base-sized BERT model on 18 English benchmark datasets related to sentiment analysis. Like Zhong et al. (2023), they found that the zero-shot performance of GPT-3.5-turbo was on par with the base-sized BERT model, and that the few-shot performance of GPT-3.5-turbo (with 27 few-shot examples) was slightly better than BERT, on average. Their test sets contained, on average, 538 samples, which is a significant improvement over Zhong et al. (2023). However, the narrow focus on the evaluation tasks as well as only benchmarking a single encoder and decoder model makes it hard to generalise the results to other tasks and models.
Kocoń et al. (2023) built a benchmark suite of 25 tasks, where 21 of these tasks were classification tasks (binary, multi-class and multi-label), 3 being question answering tasks and the last one being a token classification task. Two of the classification tasks were in Polish, the rest being English. They compared the zero-shot and few-shot performance of GPT-3.5-turbo (January 2023 version) to the state-of-the-art encoder performance on each task. GPT-3.5-turbo was generally found to be worse than state-of-the-art encoder models. They also evaluated GPT-4 on five of the tasks (inference, question-answering and emotion datasets), and only found GPT-4 to be marginally better than GPT-3.5-turbo, still far off the encoder models.
Qiu and Jin (2024) compared GPT-3.5-turbo (January 2023 version) to a finetuned version of the base-sized BERT model on three manually curated English multi-class classification datasets with 19, 12 and 7 test samples, respectively, where they found that the BERT model performed marginally better than GPT-3.5-turbo in a few-shot setting (and that the zero-shot performance was significantly worse). The tiny test sets make it hard to generalise the results, however.
2.2 Benchmarks of Generative Language Models
In recent times, several benchmarks of generative language models have been introduced. The major ones are EleutherAI’s Evaluation Harness Gao et al. (2023), Hugging Face’s Open LLM Leaderboard hug which uses the Evaluation Harness as evaluation engine, and Stanford University’s HELM Bommasani et al. (2023). These are firstly all English-only benchmarks, making it hard to generalise the results to other languages, and they only include point estimates of the dataset performance, and thus do not necessarily provide a robust assessment of the models. Further, these benchmarks are exclusively for decoder models, and thus does not provide a way to compare encoders with decoders.
There has been several language-specific benchmarks introduced as well. NorBench Samuel et al. (2023) is a collection of Norwegian evaluation datasets moreso than a dedicated evaluation framework. Further, several datasets in this collection (NorQuAD, NoReC and NorNE) are already part of ScandEval. SuperLim Berdičevskis et al. (2023) falls into the same category for Swedish. DUMB de Vries et al. (2023) is a Dutch benchmarking framework, which is only focused on encoder models. Danoliterate Holm (2024) is a Danish benchmarking framework, which has been based largely on the present ScandEval benchmark, and solely focused on evaluating decoder models. Aside from language modelling performance, the Danoliterate benchmark also measures calibration, efficiency, toxicity and fairness. While the development of language-specific benchmarks is important, it leads to too little overview of trends across benchmarks and languages and incentivises model development focused on monolingual models ignoring a potential broader appeal. ScandEval provides a unified and robust approach for comparison across model categories and Germanic languages.
Benchmarking is not the only way to evaluate language models. A new “arena approach” has been popularised by the LMSYS Arena Chiang et al. , where users can submit a prompt and get two responses from two anonymised models at random, and have to evaluate the responses. The Arena is predominantly used for English, but also currently supports six other languages. This approach is a promising way to evaluate language models, but we fear that it is not as suitable for low-resource languages due to the need of many volunteers to evaluate the responses.
Lastly, the Scandinavian Embedding Benchmark Enevoldsen et al. (2024b) complements ScandEval and focuses on evaluating embedding models on a wide range of tasks in the Scandinavian languages.
3 Datasets
In this section we present the datasets that we are evaluating the models on, all of which are now included in the ScandEval framework. We should note that these datasets either (a) already existed prior to this publication or (b) are small extensions of existing datasets. An overview of all the datasets can be found in Table 1.
3.1 Named Entity Recognition
For Norwegian, Swedish and Icelandic we use the NorNE Jørgensen et al. (2020), SUC 3.0 Gustafson-Capková and Hartmann (2006), MIM-GOLD-NER Ingólfsdóttir et al. (2020) datasets, which were already included in the ScandEval framework. For Faroese we replace the previous WikiANN-fo dataset Rahimi et al. (2019) with the new human annotated FoNE dataset Snæbjarnarson et al. (2023). We also replace the previous DaNE dataset Hvingelby et al. (2020) with the new DANSK dataset Enevoldsen et al. (2024a) covering a wider variety of domains. For German, Dutch and English we add the established NER datasets GermEval Benikova et al. (2014), the Dutch part of CoNLL-2002 Sang (2002), and the English CoNLL-2003 Sang and De Meulder (2003).
3.2 Sentiment Classification
We re-use the sentiment classification datasets AngryTweets Pauli et al. (2021), NoReC Velldal et al. (2018) and SweReC Svensson (2017), for Danish, Norwegian and Swedish, respectively. For German, Dutch and English we add the existing datasets SB10k Cieliebak et al. (2017), Dutch Social Gupta (2022) and SST5 Socher et al. (2013). We convert SST5 to the standardised trinary (negative, neutral, positive) format by converting the “very negative” and “very positive” labels to “negative” and “positive”, respectively.
3.3 Linguistic Acceptability
For linguistic acceptability we re-use the ScaLA datasets for all the Scandinavian languages, and extend the ScaLA datasets by applying the ScaLA method from Nielsen (2023) to German, Dutch and English by using the German McDonald et al. (2013), Dutch van der Beek et al. (2002) and English Zeldes (2017) dependency treebanks.
3.4 Extractive Question Answering
Here we use the ScandiQA dataset Nielsen (2023) for Danish and Swedish, but replace the manually translated Norwegian ScandiQA dataset with the new curated NorQuAD dataset Ivanova et al. (2023). We further add the new Natural Questions in Icelandic dataset Snæbjarnarson and Einarsson (2022) for Icelandic. For German and English we add the existing extractive question-answering datasets GermanQuAD Möller et al. (2021) and SQuAD Rajpurkar et al. (2016), respectively. For Dutch we add the machine translated version of SQuAD to Dutch Havinga (2023). We acknowledge that a machine translated dataset is not ideal, and will return to this point when analysing the evaluation results.
Dataset Language #Train #Val #Test #Shots
NER
      DANSK Enevoldsen et al. (2024a) Danish 1,024 256 1,024 8
      SUC 3.0 Gustafson-Capková and Hartmann (2006) Swedish 1,024 256 2,048 8
      NorNE-nb Jørgensen et al. (2020) Norwegian Bokmål 1,024 256 2,048 8
      NorNE-nn Jørgensen et al. (2020) Norwegian Nynorsk 1,024 256 2,048 8
      MIM-GOLD-NER Ingólfsdóttir et al. (2020) Icelandic 1,024 256 2,048 8
      FoNE Snæbjarnarson et al. (2023) Faroese 1,024 256 2,048 8
      GermEval Benikova et al. (2014) German 1,024 256 1,024 8
      CoNLL-nl Sang (2002) Dutch 1,024 256 1,024 8
      CoNLL-en Sang and De Meulder (2003) English 1,024 256 2,048 8
Sentiment Classification
      Angry Tweets Pauli et al. (2021) Danish 1,024 256 2,048 3
      SweReC Svensson (2017) Swedish 1,024 256 2,048 3
      NoReC Velldal et al. (2018) Norwegian 1,024 256 2,048 3
      SB10k Cieliebak et al. (2017) German 1,024 256 1,024 3
      Dutch Social Gupta (2022) Dutch 1,024 256 1,024 3
      SST5 Socher et al. (2013) English 1,024 256 2,048 3
Linguistic Acceptability
      ScaLA-da Nielsen (2023) Danish 1,024 256 2,048 3
      ScaLA-sv Nielsen (2023) Swedish 1,024 256 2,048 3
      ScaLA-nb Nielsen (2023) Norwegian Bokmål 1,024 256 2,048 3
      ScaLA-nn Nielsen (2023) Norwegian Nynorsk 1,024 256 2,048 3
      ScaLA-is Nielsen (2023) Icelandic 1,024 256 2,048 3
      ScaLA-fo Nielsen (2023) Faroese 1,024 256 1,024 3
      ScaLA-de Nielsen (2023) German 1,024 256 2,048 3
      ScaLA-nl Nielsen (2023) Dutch 1,024 256 2,048 3
      ScaLA-en Nielsen (2023) English 1,024 256 2,048 3
Question Answering
      ScandiQA-da Nielsen (2023) Danish 1,024 256 2,048 4
      ScandiQA-sv Nielsen (2023) Swedish 1,024 256 2,048 4
      NorQuAD Ivanova et al. (2023) Norwegian Bokmål 1,024 256 2,048 2
      NQiI Snæbjarnarson and Einarsson (2022) Icelandic 1,024 256 1,024 4
      GermanQuAD Möller et al. (2021) German 1,024 256 2,048 4
      SQuAD-nl Havinga (2023) Dutch 1,024 256 2,048 4
      SQuAD Rajpurkar et al. (2016) English 1,024 256 2,048 4
Table 1: All the datasets used in the NLU evaluation. Note that these have been re-sized and do not represent the sizes of the original dataset.
4 Methodology
4.1 Formulating NLU Tasks as Generative Tasks
In this section we describe how we rephrase the NLU tasks as text-to-text tasks, which makes it possible to evaluate generative models on the tasks. We formulate all the tasks as few-shot tasks, generally formatted as follows:
[prefix prompt]
[document prefix]: [document]
[label prefix]: [label]
[document prefix]: [document]
[label prefix]: [label]
[document prefix]: [document]
[label prefix]:
We found that the separation of the few-shot examples with double newlines makes it easier to know when to stop the generation - for the same reason, we ensure that there are no double newlines in any of the documents. See the prompts used for the English datasets in Table 4.1; a full table of the prompts used for all the tasks in all the languages can be found in the appendix.
For the sentiment classification task, we simply have the models generate translations of the three labels (positive, negative and neutral). For the linguistic acceptability task, also a text classification task, we use the translations of “yes” and “no” as the two labels, corresponding to whether the document is grammatically correct or not. For the extractive question answering task, we have the model output the answer directly. For this task we found that changing the label prefix from “Answer” to “Answer in max 3 words” resulted in a drastic improvement, due to many of the answers of instruction tuned models starting with unnecessary text akin to “The answer is”. Lastly, for the named entity recognition task, we require the output to be a JSON dictionary ISO/IEC 21778: (2017), with keys being the translated named entity tags, and values being lists of named entities of that category. To ensure that we are not biasing the evaluation toward models knowing the JSON format, we employ structured generation using the outlines package Louf (2023), which modifies the logits outputted by the model to ensure that the output is always a valid JSON dictionary in the aforementioned format.
{tblr}
Q[3cm,valign=m]|Q[7cm,valign=m]X[j,valign=m] Task Prefix Prompt Example Prompt
Named entity recognition Below are sentences and JSON dictionaries with the named entities that occur in the given sentence. Sentence: [text]
Named entities: [label]
Sentiment classification The following are tweets are their sentiment, which can be ’positive’, ’neutral’ or ’negative’. Tweet: [text]
Sentiment: [label]
Linguistic acceptability The following are sentences and whether they are grammatically correct. Sentence: [text]
Grammatically correct: [label]
Question answering The following are texts with accompanying questions and answers. Text: [text]
Question: [question]
Answer in max 3 words: [label]
Table 2: The English prompt templates used for the datasets. See all the prompt templates in the appendix.
4.2 Evaluation Methodology
We keep the evaluation methodology for the generative models to be as close to the methodology for encoder models in Nielsen (2023). We think of the few-shot examples as analogous to training examples for encoder models. Indeed, as von Oswald et al. (2023) shows, this assumption is theoretically grounded. We thus evaluate the models 10 times, where on each iteration we sample few-shot examples at random from the training split, and we evaluate the model on a bootstrapped version of the test split. As with the encoder models, this allows us to take into account more noise in evaluation process, resulting in more robust evaluation scores.
4.3 Score Aggregation Method
From the raw scores of the 10 evaluations per dataset, we need to aggregate the model scores into a single score. We want an aggregation method that satisfies the following criteria:
  1. 1.
    Task Fairness: Each task should be weighted equally.
  2. 2.
    Comparison: If we evaluate models in multiple languages, then it should be possible to meaningfully compare the language scores of these models with each other.
  3. 3.
    Robustness: If two models do not have a significantly different score on a dataset, then the aggregated score should reflect this.
  4. 4.
    Magnitude Preservation: The magnitude of the difference between the dataset score of two models should be reflected in the aggregated score.
  5. 5.
    Minimal Change: Adding a new model should minimally affect the aggregated scores of the other models.
Before we introduce our chosen aggregation method, we will briefly discuss some common aggregation methods and how they do not satisfy the criteria.
The mean score is the most common aggregation method, which would simply be the mean of the 10 scores for each dataset, and then the mean of the dataset scores for each task. This method does not satisfy the Task Fairness criterion, as it does not take into account that metrics have different ranges and variances. The Comparison criterion is also not satisfied, as datasets vary from language to language, with some datasets being more difficult than others. It does, however, satisfy the Robustness, Magnitude Preservation and Minimal Change criteria.
The mean rank is another common aggregation method, where we compute the rank of each model on each dataset, and then take the mean of the ranks. This method satisfies the Task Fairness criterion, as it re-casts the scores into a common comparable framework, which therefore weights each task equally. For the same reason, it also satisfies the Comparison criterion (it is important here that we evaluate all the models on all the languages for this to be satisfied). It does not satisfy the Robustness and Magnitude Preservation criteria, by definition of rank. It partially satisfies the Minimal Change criterion, since it only affects the scores of the models which are worse than the new model.
We thus see that the mean score and mean rank methods satisfy a disjoint set of the criteria, but that they together satisfy all the criteria. Based on this observation, we introduce the mean rank score method, defined as follows. For each dataset, we start by sorting the models by their mean score on the dataset. As with a rank, we assign the best model with rank score 1. For the next best model, we conduct a one-tailed Welch’s t-test to see if the next best model is significantly worse than the first model ($p<0.05$). If so, we compute the absolute difference between the mean score of the two models, and divide that by the standard deviation of all the mean scores of the models on the dataset.
We then add this to the rank score of the first model. We continue this process for all the models to get the rank scores for the dataset, and to compute the overall score for the model, we take the mean of the rank scores for the datasets. An overview of this aggregation method can be found in Algorithm 1.
We note that the mean rank score has an intuitive interpretation: it is the average number of standard deviations from the best scoring model (+1).
This metric satisfies Task Fairness since we normalise all the scores by dividing by the standard deviation of the dataset scores. The Robustness criterion is satisfied due to our use of a one-tailed Welch’s t-test. The Magnitude Preservation criterion is also satisfied, as the magnitude of the difference between the dataset score of two models is reflected in the rank score. It also satisfies Comparison, as we compare the models on a common scale (same argument as the mean rank method). Finally, the Minimal Change criterion is partially satisfied, as adding new models only minimally changes the score of existing models. Concretely, adding new scores will affect the standard deviation normalising factor (this effect tends to zero as the number of models grows, however), and if the model beats all the other models then all the scores will be affected, due to the relative nature of the metric.
Algorithm 1 The mean rank score algorithm.
1:Input: $M$ models, $D$ datasets, $S_{m,d,i}$ the $i$’th raw score of model $m$ on dataset $d$
2:Output: The mean rank score of each model
3:$R\leftarrow\textrm{matrix of size }M\times D$
4:for $\hat{d}\in D$ do
5:     $\textrm{sortedModels}\leftarrow\textrm{argsort}_{m}\nicefrac{{1}}{{10}}\cdot\Sigma_{i}S_{m,\hat{d},i}$
6:     for $\hat{m}\in\textrm{sortedModels}$ do
7:         if $\hat{m}$ is the first model then
8:              $\rho\leftarrow 1$
9:              $\textrm{previousScores}\leftarrow S_{\hat{m},\hat{d}}$
10:         else
[…]
12:              if $\textrm{pValue}<0.05$ then
13:                  $\Delta\leftarrow\textrm{mean}(S_{\hat{m},\hat{d}})-\textrm{mean}(\textrm{previousScores})$
14:                  $\rho\leftarrow\rho+\nicefrac{{\Delta}}{{\textrm{standardDeviation}_{m}S_{m,\hat{d}}}}$
15:                  $\textrm{previousScores}\leftarrow S_{\hat{m},\hat{d}}$
16:              end if
17:         end if
18:         $R_{\hat{m},\hat{d}}\leftarrow\rho$
19:     end for
20:end for
21:$\textrm{output}\leftarrow\nicefrac{{1}}{{D}}\cdot\Sigma_{d}R_{m,d}$
5 Analysis
5.1 Comparative Performance Analysis on High- and Low-resource Languages
Excerpts of the English, Danish and Icelandic leaderboards can be found in Table 3, Table 4 and Table 5, respectively. We found that these three represent three main categories of languages with respect to the open-closed source divide. Similar excerpts for the remaining languages (Swedish, Norwegian, Faroese, German and Dutch) can be found in the appendix. The full leaderboards for all the languages can be found at https://scandeval.com.
From the English results we see that the state-of-the-art decoder model GPT-4-0613 Achiam et al. (2023) is still outperformed by the DeBERTa-v3-large and DeBERTa-v3-base models He et al. (2020) as well as the ELECTRA-base model Clark et al. (2020). Here GPT-4-0613 is, on average, 0.44 standard deviations worse than the best model. The same pattern is seen for Norwegian, Dutch, German and Faroese; see the appendix for the corresponding leaderboard excerpts.
In contrast, on the Danish leaderboard, the top-3 models are all decoder models, with GPT-4-0613 and GPT-4-1106-preview OpenAI (2023b) in the lead, followed by the closed-source DanskGPT-Chat-Llama3-70B model from Syv.AI [^1] https://www.syv.ai/, being a continuation of the Llama-3-70B model AI@Meta (2024). The GPT-4-0613 model is, on average, 0.24 standard deviations from the best model. Similar results were found with Swedish; see the appendix for the corresponding leaderboard excerpt.
Lastly, for Icelandic, we see that the encoders and decoders are tied in performance, with the mDeBERTa-v3-base model and the GPT-4-1106-preview model being the top models. The GPT-4-1106-preview model is, on average, 0.24 standard deviations from the best model. We note that Icelandic is the only language where the switch from GPT-4 (gpt-4-0613) to GPT-4-turbo (gpt-4-1106-preview) resulted in a significant increase in performance. We speculate that this is due to the collaboration between OpenAI and Iceland OpenAI (2023a).
We can thus give an affirmative answer to research question (Q1), showing that decoder models can achieve significantly better NLU performance than encoder models, albeit with an order of magnitude more model parameters. For (Q3), we see that this varies between languages, but without being correlated to the language resource spectrum.
Model ID Decoder Score ($\downarrow$)
microsoft/deberta-v3-large 1.09
microsoft/deberta-v3-base 1.29
google/electra-base-discriminator 1.39
gpt-4-0613 1.44
FacebookAI/roberta-large 1.46
FacebookAI/roberta-base 1.51
microsoft/mdeberta-v3-base 1.53
gpt-4-1106-preview 1.54
gpt-4o-2024-05-13 1.64
AI-Sweden-Models/roberta-large-1160k 1.64
gpt-3.5-turbo-0613 1.78
mistralai/Mistral-7B-v0.1 1.91
Table 3: Excerpt of the English ScandEval leaderboard.
Model ID Decoder Score ($\downarrow$)
gpt-4-0613 1.24
gpt-4-1106-preview 1.25
syvai/danskgpt-chat-llama3-70b 1.29
AI-Sweden-Models/roberta-large-1160k 1.39
danish-foundation-models/encoder-large-v1 1.40
meta-llama/Meta-Llama-3-70B 1.40
AI-Sweden-Models/Llama-3-8B-instruct 1.44
gpt-4o-2024-05-13 1.46
ltg/norbert3-large 1.50
NbAiLab/nb-bert-large 1.54
vesteinn/DanskBERT 1.56
google/rembert 1.61
intfloat/multilingual-e5-large 1.62
gpt-3.5-turbo-0613 1.68
FacebookAI/xlm-roberta-large 1.71
Table 4: Excerpt of the Danish ScandEval leaderboard.
Model ID Decoder Score ($\downarrow$)
microsoft/mdeberta-v3-base 1.33
gpt-4-1106-preview 1.34
gpt-4o-2024-05-13 1.43
vesteinn/ScandiBERT-no-faroese 1.48
google/rembert 1.57
vesteinn/XLMR-ENIS 1.59
gpt-4-0613 1.79
mideind/IceBERT-large 1.85
vesteinn/FoBERT 1.87
meta-llama/Meta-Llama-3-70B 2.03
FacebookAI/xlm-roberta-large 2.34
gpt-3.5-turbo-0613 2.51
mistralai/Mistral-7B-v0.1 2.96
Table 5: Excerpt of the Icelandic ScandEval leaderboard.
5.2 Task Analysis
In this section we analyse our research question (Q2), asking whether the NLU performance results from the previous section is dependent on the type of NLU task.
Firstly, we analyse whether the score distribution across the four NLU tasks is different for the encoder and decoder models. This is done by applying a UMAP McInnes et al. (2018) to the results of a given leaderboard, stratified by both the mean rank score and whether the model is generative. UMAP plots for the English, Danish, Swedish, Norwegian, German and Dutch leaderboards can be found in Figure 1.
We see that the worst and best performing models have similar distributions, irrespective of whether they are generative or not. However, we also note that the rest of encoder and decoder models follow different “paths” in the UMAP space, leading to our hypothesis that they learn different things during their development.
Figure 1: UMAP plots of the models on the ScandEval leaderboards.
In Figure 2 we show the correlation between a model being generative and its performance on the four NLU tasks. We see that being generative is a strong predictor for good question answering performance, as well as poor named entity recognition and linguistic acceptability performance. The correlation is weaker for sentiment classification and varies across languages. We also see that these findings seem to generalise across languages, both high- and low-resource. The large question answering performance persists for non-instruction-tuned decoder models (see the leaderboards at https://scandeval.com), showing a likely side-effect of the pre-training algorithm or the architecture of decoder models making them better at this task.
Figure 2: The correlation between a model being generative and its performance on the four NLU tasks.
6 Discussion
Having a good mean rank score is not the only thing that matters when choosing a model for a given task. Model size, inference speed and whether the model has publicly available weights are all important factors to consider. For this reason, we also include these metadata in the leaderboard, and we encourage the community to consider these factors when choosing a model for a given task.
7 Conclusion
We have extended the ScandEval benchmark to include the evaluation of decoder models, as well as including three new languages: German, Dutch and English. From the analysis of the corresponding results we found that decoder models can achieve significantly better NLU performance than encoder models, but that this varies between languages. We have also shown that being generative is strongly correlated with both good question answering performance and poor performance for named entity recognition and linguistic acceptability. Our analysis showed that the “path” from the worst to the best-performing models in the UMAP space is different for encoder and decoder models, indicating that they learn different things during their development.
Ethics Statement
We have taken care to ensure that the evaluation is fair and unbiased, both through our selection of the datasets in the benchmark as well as through our choice of aggregation method of the scores.
Acknowledgements
This work has received funding by the European Union’s Horizon 2023 Research and Innovation Actions, as part of the Artificial Intelligence and Robotics programme, for the project “TrustLLM” (grant agreement number 101135671). Furthermore, this work reflects only the authors’ view and the European Research Executive Agency (REA) is not responsible for any use that may be made of the information it contains.
Appendix A Appendix
Model ID Decoder Score ($\downarrow$)
AI-Sweden-Models/Llama-3-8B-instruct 1.07
gpt-4-0613 1.28
four-two-labs/lynx-micro 1.29
AI-Sweden-Models/roberta-large-1350k 1.33
meta-llama/Meta-Llama-3-70B 1.34
gpt-4o-2024-05-13 1.35
gpt-4-1106-preview 1.37
KBLab/megatron-bert-large-swedish-cased-165k 1.42
AI-Sweden-Models/bert-large-nordic-pile-1M-steps 1.48
KB/bert-base-swedish-cased 1.48
intfloat/multilingual-e5-large 1.49
ltg/norbert3-large 1.49
gpt-3.5-turbo-0613 1.57
upstage/SOLAR-10.7B-v1.0 1.96
timpal0l/Mistral-7B-v0.1-flashback-v2 2.02
mistralai/Mistral-7B-v0.1 2.17
Table 6: Excerpt of the Swedish ScandEval leaderboard.
Model ID Decoder Score ($\downarrow$)
ltg/norbert3-large 1.12
NbAiLab/nb-bert-large 1.19
ltg/norbert3-base 1.32
AI-Sweden-Models/roberta-large-1350k 1.38
gpt-4-0613 1.49
danish-foundation-models/encoder-large-v1 1.54
google/rembert 1.57
meta-llama/Meta-Llama-3-70B 1.59
gpt-4-1106-preview 1.62
gpt-4o-2024-05-13 1.67
microsoft/mdeberta-v3-base 1.67
gpt-3.5-turbo-0613 2.03
upstage/SOLAR-10.7B-v1.0 2.38
meta-llama/Meta-Llama-3-8B 2.45
mistralai/Mistral-7B-v0.1 2.79
AI-Sweden-Models/Llama-3-8B-instruct 2.82
Table 7: Excerpt of the Norwegian ScandEval leaderboard.
Model ID Decoder Score ($\downarrow$)
vesteinn/FoBERT 1.00
microsoft/mdeberta-v3-base 2.01
vesteinn/ScandiBERT-no-faroese 2.66
gpt-4-1106-preview 2.69
google/rembert 3.35
ltg/norbert3-base 3.36
gpt-4-0613 3.46
mideind/IceBERT-large 3.96
meta-llama/Meta-Llama-3-70B 4.16
gpt-3.5-turbo-0613 4.23
gpt-4o-2024-05-13 4.69
mistralai/Mistral-7B-v0.1 4.84
Table 8: Excerpt of the Faroese ScandEval leaderboard.
Model ID Decoder Score ($\downarrow$)
deepset/gbert-large 1.17
FacebookAI/xlm-roberta-large 1.34
google/rembert 1.34
gpt-4-0613 1.49
intfloat/multilingual-e5-large 1.49
meta-llama/Meta-Llama-3-70B 1.57
gwlms/deberta-base-dewiki-v1 1.61
gpt-4-1106-preview 1.65
upstage/SOLAR-10.7B-v1.0 1.67
microsoft/mdeberta-v3-base 1.74
gpt-4o-2024-05-13 1.82
deepset/gbert-base 1.83
gpt-3.5-turbo-0613 1.95
occiglot/occiglot-7b-de-en-instruct 2.02
meta-llama/Meta-Llama-3-8B 2.02
mistralai/Mistral-7B-v0.1 2.15
Table 9: Excerpt of the German ScandEval leaderboard.
Model ID Decoder Score ($\downarrow$)
intfloat/multilingual-e5-large 1.42
gpt-4-0613 1.69
DTAI-KULeuven/robbert-2022-dutch-base 1.78
pdelobelle/robbert-v2-dutch-base 1.81
meta-llama/Meta-Llama-3-70B 1.84
gpt-4-1106-preview 1.98
intfloat/multilingual-e5-base 2.00
FacebookAI/xlm-roberta-large 2.10
gpt-4o-2024-05-13 2.14
microsoft/mdeberta-v3-base 2.25
gpt-3.5-turbo-0613 2.27
google/rembert 2.29
upstage/SOLAR-10.7B-v1.0 2.33
meta-llama/Meta-Llama-3-8B 2.54
mistralai/Mistral-7B-v0.1 2.77
Table 10: Excerpt of the Dutch ScandEval leaderboard.
{tblr}
Q[3cm,valign=m]|Q[7cm,valign=m]X[j,valign=m] Dataset Prefix Prompt Example Prompt
DANSK Følgende er sætninger og JSON-ordbøger med de navngivne enheder, som forekommer i den givne sætning. Sætning: [text]
Navngivne enheder: [label]
SUC 3.0 Följande är meningar och JSON-ordböcker med de namngivna enheter som förekommer i den givna meningen. Mening: [text]
Namngivna entiteter: [label]
NorNE-nb
NorNE-nn Følgende er fraser og JSON-ordbøker med de navngitte enhetene som forekommer i den gitte frasen. Frase: [text]
Navngitte enheter: [label]
MIM-GOLD-NER Eftirfarandi eru setningar ásamt JSON lyklum með nefndum einingum sem koma fyrir í setningunum. Setning: [text]
Nefndar einingar: [label]
FoNE Her eru nakrir setningar og nakrar JSON orðabøkur við nevndar eindir, sum eru í setningunum. Setningur: [text]
Nevndar eindir: [label]
GermEval Es folgen Sätze und JSON-Wörterbücher mit den benannten Entitäten, die in der angegebenen Phrase vorkommen. Satz: [text]
Benannte Entitäten: [label]
CoNLL-nl Hieronder staan zinnen en JSON woordenboeken met de genoemde entiteiten die voorkomen in de gegeven zin. Zin: [text]
Genoemde entiteiten: [label]
CoNLL-en Below are sentences and JSON dictionaries with the named entities that occur in the given sentence. Sentence: [text]
Named entities: [label]
Table 11: The prompt templates used for the NER datasets.
{tblr}
l|Q[7cm,valign=m]X[j,valign=m] Dataset Prefix Prompt Example Prompt
Angry Tweets Følgende er tweets og deres sentiment, som kan være ’positiv’, ’neutral’ eller ’negativ’. Tweet: [text]
Sentiment: [label]
SweReC Följande är recensioner och deras sentiment, som kan vara ’positiv’, ’neutral’ eller ’negativ’. Recension: [text]
Sentiment: [label]
NoReC Følgende er anmeldelser og deres sentiment, som kan være ’positiv’, ’nøytral’ eller ’negativ’. Anmeldelse: [text]
Sentiment: [label]
SB10k Im Folgenden sind Tweets und ihre Stimmung aufgeführt, die ’positiv’, ’neutral’ oder ’negativ’ sein kann. Tweet: [text]
Stimmungslage: [label]
Dutch Social Hieronder staan tweets en hun sentiment, dat ’positief’, ’neutraal’ of ’negatief’ kan zijn. Tweet: [text]
Sentiment: [label]
SST5 The following are tweets are their sentiment, which can be ’positive’, ’neutral’ or ’negative’. Tweet: [text]
Sentiment: [label]
Table 12: The prompt templates used for the sentiment classification datasets.
{tblr}
Q[2cm,valign=m]|Q[7cm,valign=m]X[j,valign=m] Dataset Prefix Prompt Example Prompt
ScaLA-da Følgende er sætninger og om de er grammatisk korrekte. Sætning: [text]
Grammatisk korrekt: [label]
ScaLA-sv Följande är meningar och huruvida de är grammatiskt korrekta. Mening: [text]
Grammatisk korrekt: [label]
ScaLA-nb
ScaLA-nn Følgende er setninger og hvorvidt de er grammatisk korrekte. Setning: [text]
Grammatisk korrekt: [label]
ScaLA-is Eftirfarandi eru setningar og hvort þær eru málfræðilega réttar. Setning: [text]
Málfræðilega rétt: [label]
ScaLA-fo Hetta eru nakrir setningar og um teir eru mállæruliga rættir. Setningur: [text]
Mállæruliga rættur: [label]
ScaLA-de Die folgenden Sätze und ob sie grammatikalisch korrekt sind. Satz: [text]
Grammatikalisch richtig: [label]
ScaLA-nl Hieronder staan zinnen en of ze grammaticaal correct zijn. Zin: [text]
Grammaticaal correct: [label]
ScaLA-en The following are sentences and whether they are grammatically correct. Sentence: [text]
Grammatically correct: [label]
Table 13: The prompt templates used for the linguistic acceptability datasets.
{tblr}
l|Q[5cm,valign=m]X[j,valign=m] Dataset Prefix Prompt Example Prompt
ScandiQA-da Følgende er tekster med tilhørende spørgsmål og svar. Tekst: [text]
Spørgsmål: [question]
Svar med maks. 3 ord: [label]
ScandiQA-sv Her følger tekster med tilhørende spørsmål og svar. Tekst: [text]
Spørsmål: [question]
Svar på maks 3 ord: [label]
NorQuAD Nedan följer texter med tillhörande frågor och svar. Text: [text]
Fråga: [question]
Svar på max 3 ord: [label]
NQiI Eftirfarandi eru textar með tilheyrandi spurningum og svörum. Texti: [text]
Spurning: [question]
Svaraðu með að hámarki 3 orðum: [label]
GermanQuAD Im Folgenden finden Sie Texte mit den dazugehörigen Fragen und Antworten. Text: [text]
Fragen: [question]
Fragen Antwort in maximal 3 Wörtern: [label]
SQuAD-nl Hieronder volgen teksten met bijbehorende vragen en antwoorden. Tekst: [text]
Vraag: [question]
Antwoord in max 3 woorden: [label]
SQuAD The following are texts with accompanying questions and answers. Text: [text]
Question: [question]
Answer in max 3 words: [label]
Table 14: The prompt templates used for the question answering datasets.
{tblr}
l|cccccccc Language PER LOC ORG Negative Neutral Positive Correct Incorrect
Danish Person Sted Organisation Negativ Neutral Positiv Ja Nej
Swedish Person Plats Organisation Negativ Neutral Positiv Ja Nej
Norwegian Person Sted Organisasjon Negativ Nøytral Positiv Ja Nei
Icelandic Einstaklingur Staðsetning Stofnun Já Nei
Faroese Persónur Staður Felagsskapur Ja Nei
German Person Ort Organisation Positiv Neutral Negativ Ja Nein
Dutch Persoon Locatie Organisatie Positief Neutraal Negatief Ja Nee
English Person Location Organization Negative Neutral Positive Yes No
Table 15: The label conversions for the datasets.